OpenStudy (anonymous):

i need the steps to multiplying 26/2 x 5 i know the answer but i need to know how i got it.

OpenStudy (anonymous):

u divide 26 by 2 to get 13 and then mulitply it by 5 to get 65 ??

OpenStudy (anonymous):

i did that on one of my other problems and didnt get the right answer which is why i assumed that i was doing it wrong. thanks!

OpenStudy (anonymous):

is that the answer u wanted ? or have u got something else?

OpenStudy (anonymous):

the answer:5i

OpenStudy (anonymous):

the answer:5i

OpenStudy (anonymous):

sorry : 65i

OpenStudy (anonymous):

but there is no i(iota or complex no.) in ur problem ?

OpenStudy (anonymous):

26/2=13 (13)*(5i)=65i

OpenStudy (anonymous):

yes thats right.....